3(x-3)+4(x+1)= -12
Answer by checkley77(12844)

3(x-3)+4(x+1)= -12
3X-9+4X+4=-12
7X=-12+9-4
7X=-7
X=-7/7
X=-1 ANS.
PROOF:
3(-1-3)+4(-1+1)=12
3*-4+0=-12
-12=-12